OpenStudy (anonymous):

The volume of a cylinder is 81 ft3. The volume of a cube is 42 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube?

OpenStudy (anonymous):

V=pir^2*h cylinder v=1/3pir^2h cone v=l*w*h V=1/3b^2h 1/3 of the volume can fit a whole shape of a cone

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@Luigi0210

jimthompson5910 (jim_thompson5910):

The volume of a cylinder is 81 ft3 so what is the volume of the cone that can fit inside it?

OpenStudy (anonymous):

39

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I really don't understand how to do this please explain it to me

jimthompson5910 (jim_thompson5910):

Volume of Cone = (1/3)*(Volume of Cylinder) This is only if the cone and cylinder have the same radius and height

OpenStudy (anonymous):

So how do we know if they have the same radius and height

jimthompson5910 (jim_thompson5910):

because the smaller figure fits inside the larger one

OpenStudy (anonymous):

Ok so how do we write the equation 81 ft3/42 ft3. volume of Cone = (1/3)*(Volume of Cylinder)

jimthompson5910 (jim_thompson5910):

what is the Volume of Cylinder

OpenStudy (anonymous):

81 ft3 can you write the equation I'll work it out

jimthompson5910 (jim_thompson5910):

the equation is volume of Cone = (1/3)*(Volume of Cylinder)

jimthompson5910 (jim_thompson5910):

plug in the given number and evaluate to find the volume of cone

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

similarly, volume of pyramid = (1/3)*(volume of cube)

OpenStudy (anonymous):

14

jimthompson5910 (jim_thompson5910):

now find the difference between the cone and the pyramid

jimthompson5910 (jim_thompson5910):

the volumes of each

OpenStudy (anonymous):

81-21=60

jimthompson5910 (jim_thompson5910):

not the cylinder and cube

jimthompson5910 (jim_thompson5910):

the cone and the pyramid

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

and idk where you got 21 from

OpenStudy (anonymous):

21-14=7

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I am lost

jimthompson5910 (jim_thompson5910):

go back up and see what the volume of the cone and pyramid were

OpenStudy (anonymous):

The cone is larger than the pyramid.

jimthompson5910 (jim_thompson5910):

what were the volumes of each

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

that's the cone's volume

OpenStudy (anonymous):

14

OpenStudy (anonymous):

27-14=13

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

sorry about the 21 i hit one instead of 8

OpenStudy (anonymous):

thank for all you help i really didn't know how to do it

jimthompson5910 (jim_thompson5910):

it's ok

OpenStudy (anonymous):

i mint 7 not 8